# Hierarchical Clustering Algorithms on Poisson and Cox Point Processes

### v2

@@ 2.1 Construction of the Pre-limit Point-Shifts
One defines $f^{k}$ from $f^{k-1}$ by keeping the images of all points the same except for exit points of order $k-1$. The image of such a point say $x\in S^{k-1}_{i}$ define $f^{k}(x)=\mathrm{nnk}(x)$.
Each connected component of the graph of $f^{k}$ on $\Phi^{0}$ is referred to as a cluster of order k. It will be shown in Corollary 5.7 that these connected components are almost surely $\mathcal{F/F}$, meaning that each cluster in the graph of $f^{k}$ is finite and contains a unique cycle. These cycles are of length 2 a.s. and connect two exit points of order $k-1$, denoted $x$ and $y$, and such that $\mathrm{nnk}(x)=y$ and $\mathrm{nnk}(y)=x$. These cycles are called $k$-cycles.
@@
Figure 7: Left picture: the graph of $f^{3}$ with the $3$-cycles in boldface;
Right picture: Cluster subtrees of order 3.
Remark 2.6.
@@ 2.2 Construction of the Limiting Point-Shift
Corollary 2.8 (Construction of the point-shift graph).
@@
This is assessed under the Palm probability of $\Phi^{0}$. For all balls of radius $h$ centered at 0, the restriction of $f^{n}$-graphs to this ball converges a.s. to a limit when $n$ tends to infinity. This is because the intensity of the points of $\Phi^{i}_{c}$ tends to 0 as $i$ tends to infinity. Hence, there is a random $n$ such that the restriction of the graph $G^{n}$ to this ball is fully determined by $f^{n}$ (this random integer is the largest integer $n$ such that the ball in question is empty of points of $\Phi^{n+1}_{c}$). We will denote by $[G,0]$ the limiting random graph, and by $[G^{n},0]$ the $f^{n}$-graph under the Palm probability of $\Phi^{0}$. Since for each $n$, $[G^{n},0]$ is a unimodular graph (see [5]) the limit graph $[G,0]$ is also unimodular (see [6]). ∎
@@
For each $n\in\mathbb{N}$, the connected components of the graph $[G^{n},0]$ are all of type $\mathcal{F/F}$.
@@
In Section 5.1, it will be shown that, for all $n$, there is no infinite path in $[G^{n},0]$. This fact demonstrates that, for each $n$, all the clusters of $[G^{n},0]$ are finite. Note that the cycle indicated in the properties of the $\mathcal{F/F}$ components (see Theorem 3.10 in [6]) in each cluster of $f^{n}$-graph is the cycle of mutual nearest neighbor points, which is a cycle of length two a.s. ∎
Remark 2.10.
@@
It is known that there exists a directed graph $G^{\prime}$ with vertex set $\Phi^{0}$, that is constructed in a deterministic and isometry-invariant way, which is a.s. a doubly infinite path (see [17]). Consequently, a point-shift on $\Phi^{0}$ can be obtained by mapping each vertex to its subsequent vertex following the direction of this path. This point-shift is a bijection. By applying Mecke’s Point Stationary Theorem (see [5]), it can be concluded that the distribution of $[G,0]$ is identical to the distribution of $[G,v]$. ∎
@@
The connected components of $\mathrm{CHN}^{2}$ EFF $[G,0]$ belong to the $\mathcal{I/I}$ class of the foil classification theorem for unimodular networks. In other words, under the Poisson Point Process (PPP), each connected component of $\mathrm{CHN}^{2}$ EFF is a one-ended tree with all its foils infinite almost surely.
@@
Remark 2.13 (Evaporation and $\mathcal{I/I}$ Property).
Corollary 2.12 can also be proved using the evaporation property of the $\mathrm{CHN}^{2}$ point-shift, $f^{\infty}$, on the $\mathrm{CHN}^{2}$ EFF, $[G,0]$. Recall that a point-shift $f$ on a connected component $C\subseteq[G,0]$ evaporates $C$ if, almost surely, there is no point $x\in C$ with arbitrarily large backward iterates under $f$, i.e., $f^{-k}(x)=\emptyset$ almost surely for all sufficiently large $k$ (see [5] for more details). Corollary 24 in [5] shows that if a point-shift on a point process evaporates a connected component $C$ of its point-shift graph, then $C$ belongs to the $\mathcal{I/I}$ class. Since the intensity of exit points at level $k$ goes to zero as $k\to\infty$, it follows that $f^{\infty}$ evaporates every connected component of $[G,0]$. Hence, all such components belong to the $\mathcal{I/I}$ class.
@@
Theorem 2.14.
In dimension $d=1$, under the Palm probability of $\Phi^{0}\subseteq\mathbb{R}$, the $\mathrm{CHN}^{2}$ EFF is a tree.
@@
To prove that $[G,0]$ contains only one connected component in dimension $1$, it suffices to show that for every fixed $r>0$, almost surely all points of $\Phi^{0}$ within the ball of radius $r$ centered at 00, denoted as $B_{r}(0)$, belong to the same component.
Recall that at each level $n$ of the $\mathrm{CHN}^{2}$ clustering algorithm, the point process $\Phi^{n}_{c}$ consists of the vertices of $f^{n}$-cycles $S_{i}^{n}$ at level $n$, (see definition in Subsection 2.1). Note that the distribution of distances between pairs of $S_{i}^{n}$ cycles, for each $i$, is stationary. Consider the following events:
$A_{n+1}=\{\text{The first $f^{n}$-cycle after point $0$ in the positive side of $\mathbb{R}$ is connected}$
$\text{to an $f^{n}$-cycle in the negative side of $\mathbb{R}$ atlevel $n+1$ of the algorithm}\}.$
Consider the two points, $x_{i}$ and $y_{i}$, which are part of the $f^{n}$-cycles belonging to $S^{n}_{i}$. Let’s define $M_{i}^{n}=\frac{x_{i}+y_{i}}{2}$, which represents the middle points of these two points. The sequence of random distances between $M_{i}^{n}$s, $\{d_{i}\}_{i\in\mathbb{Z}}=\{|M_{i}^{n}-M_{i-1}^{n}|\}_{i\in\mathbb{Z}}$, forms a stationary process. Therefore, the probability of a typical cycle connecting to its left neighbor cycle is equal to the probability of it connecting to its right neighbor cycle. Therefore,
$\sum_{n}\mathbb{P}(A_{n})=\infty.$
The Borel-Cantelli lemma implies that $A_{n}$ occurs for infinitely many $n$. Similarly, one can show that the same result holds for each n, for the event $A^{\prime}_{n}$,
$A^{\prime}_{n+1}=\{\text{The first $f^{n}$-cycle after point $0$in the negative side of $\mathbb{R}$ is connected}$
$\text{to an $f^{n}$-cycle in the positive side of $\mathbb{R}$ atlevel $n+1$ of the algorithm}\}.$
Furthermore, note that when a cluster from the right-hand side of zero connects to a cluster on the left-hand side of zero and forms a new cluster, all points between them also belong to the same cluster. It is also evident that at step $n$, the place of the first $f^{n}$-cycle from the right-hand side of zero tends to infinity as $n\to\infty$. Thus, for each $r$, there exists a sufficiently large $n$ such that the first $f^{n}$-cycle from the right-hand side of zero connects to the left-hand side, and $B_{r}(0)$ is contained in the new cluster formed by connecting these $f^{n}$-cycles to each other. ∎
@@
Remark 2.15.
All the results shown here for the PPP case are due to its property of not admitting a second-order descending chain (see Definition 5.1). Therefore, for any stationary point process that satisfies this property, the same results hold.
@@ 2.2.2 The Cardinality of Descendant Trees
More precisely, let the intensity of the cluster heads of level $k$, $\Phi_{c}^{k}$, be $\rho_{k}$. Then the mean of the cardinality of the descendant tree of a point belongs to $\Phi_{c}^{k}$, is $1/\rho_{k}$. It is known from Remark 2.6 that $\rho_{k}$ tend to zero, as $k$ goes to infinity.
@@
Let $G$ be the $\mathrm{CHN}^{2}$ EFF under the Palm probability of $\Phi^{0}\subseteq\mathbb{R}^{d}$. Let $N^{0}$ be the cardinality of the descendant tree of 00. Then the mean of $N^{0}$ is infinite. So $N^{0}$ is heavy-tailed.
@@
The rooted graph $[G,0]$ is a unimodular graph. Let $f$ be the $\mathrm{CHN}^{2}$ point-shift on $[G,0]$. Define $g[G,x,y]=\sum_{n=1}^{\infty}\mathbf{1}_{y=f^{n}(x)}$, which is well-defined and measurable. Therefore, the mass transport principle (2) holds for $g$ (see [2]).
$\mathbb{E}[\sum_{v\in V(\mathbf{G})}g[\mathbf{G},o,v]]=\mathbb{E}[\sum_{v\in V(\mathbf{G})}g[\mathbf{G},v,o]].$ (2)
The left-hand side of (2) for $g$ is equal to infinity, and its right-hand side is equal to the mean number of descendants of 00. So the mean of $N^{0}$ for $[G,0]$ is infinite. ∎
Remark 2.17.
@@
3 Single Linkage and Hausdorff Hierarchical Nearest Neighbor Clustering Algorithm
In the $\mathrm{CHN}^{2}$ clustering algorithm defined in Sections 1 and 2, the distance between clusters is based on the centroid distance. However, alternative definitions for the distance between clusters can also be used, such as the single linkage distance which is defined for all $X,Y\subseteq\mathbb{R}^{d}$ as follows:
$d_{S}(X,Y):=\inf\left\{d(x,y):x\in X,y\in Y\right\},$ (3)
or the Hausdorff (complete linkage) distance, which is defined for all $X,Y\subseteq\mathbb{R}^{d}$ as follows:
$d_{H}(X,Y):=\max\left\{\sup_{x\in X}d(x,Y),\sup_{y\in Y}d(X,y)\right\},$ (4)
where $d(a,B):=\inf_{b\in B}d(a,b)$ quantifies the distance from a point $a$ to a subset $B$.
While using these alternative distances in the algorithm requires more processing steps than in the $\mathrm{CHN}^{2}$ algorithm, it is still possible to define each level of the algorithm in a similar way, except for the fact that only undirected graphs will be constructed here. These algorithms will only be applied to PPP.
Let $D\subset\mathbb{R}^{d}$ be a dataset. The definitions of the Single Linkage Hierarchical Nearest Neighbor ($\mathrm{SHN}^{2}$) and the Hausdorff (Complete Linkage) Hierarchical Nearest Neighbor ($\mathrm{H}^{2}\mathrm{N}^{2}$) clustering algorithms are as follows:
Level 0 The zero level is defined similarly to the $\mathrm{CHN}^{2}$ clustering algorithm. Each point in $D$ is connected to its nearest neighbor with an undirected edge, resulting in a collection of finite connected components (Theorem 2.1 for $n=0$) called clusters of order 0. In order to ensure that the resulting graphs do not contain cycles, only one edge of each mutual nearest neighbor pair is retained. Consequently, the resulting graph consists of clusters that form undirected trees (see Figure 8).
Figure 8: Clusters of order zero generated by the $\mathrm{SHN}^{2}$ (and $\mathrm{H}^{2}\mathrm{N}^{2}$) clustering algorithm on a finite dataset. The blue edges connect each point to its nearest neighbor. A single red edge connects points that are mutual nearest neighbors. Note that the resulting graph forms an undirected forest.
Level 1 Each cluster of order 0 is connected to its nearest neighboring cluster based on the single linkage distance between clusters for the $\mathrm{SHN}^{2}$ algorithm and the Hausdorff distance for the $\mathrm{H}^{2}\mathrm{N}^{2}$ algorithm. Specifically:
  • For the $\mathrm{SHN}^{2}$ algorithm, define the proximity point of a cluster of order 0 as the point of the cluster that is closest to its nearest neighbor cluster based on the single linkage distance (3). Connect the proximity point to its nearest neighbor in the nearest neighbor cluster.
  • For the $\mathrm{H}^{2}\mathrm{N}^{2}$ algorithm, define the proximity point as the point within the cluster that attains the supremum in the Hausdorff distance (4) with respect to its nearest neighboring cluster. Connect the two points that attain the supremum in the Hausdorff distance (4).
This defines clusters of level 1. Since the finiteness of clusters at this level is not established here, they may be either finite or infinite. If a cluster is finite, it contains a mutual nearest neighbor pair. To ensure that the resulting graphs do not contain cycles, only one edge of each mutual nearest neighbor pair is retained. Consequently, the resulting graph consists of clusters that form undirected trees (see Figure 9).
Figure 9: Clusters of order one produced by the $\mathrm{SHN}^{2}$ clustering algorithm on a finite dataset. Each order-0 cluster is connected to its nearest neighbor cluster using the single linkage distance. The green edges indicate the new connections between clusters, while the pink points highlight the proximity points of order 0, that is, the points in each cluster that are closest to their respective nearest neighbor cluster. Note that the resulting graph forms an undirected forest.
Level $i$ ($i\geq 2$): At this stage, it is not guaranteed that the clusters of order 1 and higher are almost surely finite. Therefore, the algorithm is generalized to handle the possibility of infinite clusters, in order to extend the procedure to higher levels. At level $i$, each finite cluster of order $i-1$ is connected to its nearest neighbor cluster using either the single linkage or the Hausdorff distance, depending on the algorithm. Consequently, the notion of proximity points applies only to finite clusters at each level (see Figure 10).
In the $\mathrm{H}^{2}\mathrm{N}^{2}$ algorithm, if the nearest neighbor of a finite cluster $C_{j}$ is the infinite cluster $C_{j}^{\prime}$ and the maximum defined in (4) is infinite, we proceed as follows: instead of directly relying on the Hausdorff distance, we identify two vertices, one in $C_{j}$ and one in $C_{j}^{\prime}$, that achieve the infimum in (3), which corresponds to the single linkage distance between the two clusters. We then connect these two vertices, and the vertex in $C_{j}$ is defined as its proximity point. By iterating this procedure, one can proceed hierarchically as above to define proximity points of order 1 or higher and clusters of order 2 or higher.
Figure 10: Clusters of order two produced by the $\mathrm{SHN}^{2}$ clustering algorithm on a finite dataset. Each finite order-1 cluster is connected to its nearest neighbor cluster using the single linkage distance. The orange edges indicate the new connections between clusters, while the cyan points highlight the proximity points of order 1, that is, the points in each order-1 cluster that are closest to their respective nearest neighbor cluster. Note that the resulting graph again forms an undirected forest.
In the following, we analyze the graphs obtained by considering these algorithms on a PPP.
Definition 3.1.
Let $\Phi^{0}$ be a Poisson point process with intensity 1, and $[\mathcal{G}^{n}_{\mathcal{{S}}},0]$ (resp. $[\mathcal{G}^{n}_{\mathcal{{H}}},0]$) be the random undirected rooted graph, rooted at zero, constructed at the $n$-th level of the $\mathrm{SHN}^{2}$($\mathrm{H}^{2}\mathrm{N}^{2}$) clustering algorithm under the Palm probability of $\Phi^{0}$. The vertices of $[\mathcal{G}^{n}_{\mathcal{{S}}},0]$ (resp. $[\mathcal{G}^{n}_{\mathcal{{H}}},0]$) that are proximity points of level $n$ form a sub-PP of $\Phi^{0}$ denoted by $\Xi_{S}^{n}$ (resp. $\Xi_{H}^{n}$) for the $\mathrm{SHN}^{2}$ (resp. $\mathrm{H}^{2}\mathrm{N}^{2}$) clustering algorithm.
The following theorem shows that the weak limit graph of the graphs $[\mathcal{G}^{n}_{\mathcal{{S}}},0]$ exists as $n\to\infty$.
Theorem 3.2.
The weak limit graph of $[\mathcal{G}^{n}_{\mathcal{{S}}},0]$ (and similarly $[\mathcal{G}^{n}_{\mathcal{{H}}},0]$) in Definition 3.1 exists as $n\to\infty$.
Proof.
See Subsection 5.2. ∎
Definition 3.3.
The weak limit graph of the random graphs $[\mathcal{G}^{k}_{\mathcal{{S}}},0]$ in Definition 3.1 is called the Single Linkage Hierarchical Nearest Neighbor ($\mathrm{SHN}^{2}$) Spanning Forest and is denoted by $[\mathcal{G}_{\mathcal{{S}}},0]$. If this forest happens to be a single connected component (i.e., a tree), it is referred to as the Single Linkage Hierarchical Nearest Neighbor ($\mathrm{SHN}^{2}$) Spanning Tree.
There is a close relation between the Minimal Spanning Forest (MSF) and the $\mathrm{SHN}^{2}$ Spanning Forest on a point process in $\mathbb{R}^{d}$ for $d\geq 2$. The Minimal Spanning Forest, or more precisely the Wired MSF on the point process $\Phi$, was defined by Aldous and Steele in [23] and [1] as the union of all invasion percolation trees started at every vertex in $\Phi$. Invasion percolation from $x$ grows a tree by iteratively adding the shortest available edge that connects the existing tree to a new point (see also [19]). The following corollary describes this relationship.
Corollary 3.4.
Let $\Phi^{0}$ be a Poisson point process on $\mathbb{R}^{d}$ with $d\geq 2$. The $\mathrm{SHN}^{2}$ Spanning Forest is a subgraph of the MSF on $\Phi^{0}$.
Proof.
Under the Palm probability of $\Phi^{0}$, consider a fixed realization $\phi_{0}$. It suffices to show that every edge in $[\mathcal{G}^{n}_{\mathcal{{S}}},0]$ (the $n$-th step of the $\mathrm{SHN}^{2}$ clustering algorithm) also appears in the Minimal Spanning Forest (MSF) of $\phi_{0}$.
Base Case $n=0$. At step zero, the $\mathrm{SHN}^{2}$ algorithm produces the nearest neighbor graph, where each point in $\phi_{0}$ connects to its nearest neighbor. Such edges are part of any spanning tree generated by invasion percolation, and thus belong to the MSF. Hence the statement holds for $n=0$.
Inductive Step. Assume that for some $k\in\mathbb{N}$, all edges in $[\mathcal{G}^{k}_{\mathcal{{S}}},0]$ are contained in the MSF of $\phi_{0}$. We need to show this also holds for $[\mathcal{G}^{k+1}_{\mathcal{{S}}},0]$.
Let $e_{xy}$ be an edge that appears for the first time at step $k+1$. By definition of the $\mathrm{SHN}^{2}$ algorithm, $e_{xy}$ connects two clusters $C_{x}$ and $C_{y}$ built at step $k$. Assume, without loss of generality, that $C_{y}$ is the nearest neighbor of $C_{x}$. By definition, $C_{x}$ is finite.
For the sake of reaching a contradiction, suppose that $e_{xy}$ is not in the MSF of $\phi_{0}$. Consider the invasion percolation tree $t_{x}^{\infty}$ rooted at $x$. This tree is infinite, but does not include the edge $e_{xy}$. A first observation is that $C_{x}$ and $t_{x}^{\infty}$ have a non-empty intersection (both contain $x$). Since $t_{x}^{\infty}$ is infinite, whereas $C_{x}$ is finite, $t_{x}^{\infty}$ contains at least one edge that connects $C_{x}$ to a point (resp. a cluster of order $k$) that does belong to $C_{x}$ (resp. different from $C_{x}$). Every such edge of $t_{x}^{\infty}$ must have length strictly less than $\lvert e_{xy}\rvert$. But this contradicts the assumption that $C_{y}$ is the nearest neighboring cluster of $C_{x}$. Thus, $e_{xy}$ must be part of the MSF.
By induction, all new edges introduced at each level $n$ also lie in the MSF. Hence $[\mathcal{G}^{n}_{\mathcal{{S}}},0]$ is a subgraph of the MSF for every $n$. Since the weak limit of $[\mathcal{G}^{n}_{\mathcal{{S}}},0]$ exists and defines the $\mathrm{SHN}^{2}$ Spanning Forest on $\Phi^{0}$, it follows that the $\mathrm{SHN}^{2}$ Spanning Forest is a subgraph of the MSF on $\Phi^{0}$. ∎
Remark 3.5.
The definition of the $\mathrm{SHN}^{2}$ Spanning Forest is local, and studying its properties may provide insights into the behavior of the MSF on PPP. For example, if the $\mathrm{SHN}^{2}$ Spanning Forest is connected on a PPP in some dimension, then the MSF is also connected in that dimension.
@@ 3 Single Linkage and Hausdorff Hierarchical Nearest Neighbor Clustering Algorithm
4 Hierarchical Clustroid Nearest Neighbor Chain Algorithm on Cox Point Processes

### v3

@@ 2.1 Construction of the Pre-limit Point-Shifts
One defines $f^{k}$ from $f^{k-1}$ by keeping the images of all points the same except for exit points of order k1k-1. The image of such a point say $x\in S^{k-1}_{i}$ define $f^{k}(x)=\mathrm{nnk}(x)$.
Each connected component of the graph of $f^{k}$ on $\Phi^{0}$ is referred to as a cluster of order k. It will be shown in Corollary 4.6 that these connected components are almost surely $\mathcal{F/F}$, meaning that each cluster in the graph of $f^{k}$ is finite and contains a unique cycle. These cycles are of length 2 a.s. and connect two exit points of order k1k-1, denoted $x$ and $y$, and such that $\mathrm{nnk}(x)=y$ and $\mathrm{nnk}(y)=x$. These cycles are called $k$-cycles.
@@
Remark 1.
@@ 2.2 Construction of the Limiting Point-Shift
Theorem 2.7 (Construction of the point-shift graph).
@@
This is assessed under the Palm probability of $\Phi^{0}$. For all balls of radius $h$ centered at 0, the restriction of $f^{n}$-graphs to this ball converges a.s. to a limit when $n$ tends to infinity. This is because the intensity of the points of $\Phi^{i}_{c}$ tends to 0 as $i$ tends to infinity. Hence, there is a random $n$ such that the restriction of the graph $G^{n}$ to this ball is fully determined by $f^{n}$ (this random integer is the largest integer $n$ such that the ball in question is empty of points of $\Phi^{n+1}_{c}$). We will denote by $[G,0]$ the limiting random graph, and by $G^{n},0]$ the $f^{n}$-graph under the Palm probability of $\Phi^{0}$. Since for each $n$, $G^{n},0]$ is a unimodular graph (see [haji_mirsadeghi_baccelli2018]) the limit graph $[G,0]$ is also unimodular (see [Baccelli2017]). ∎
@@
For each $n\in\mathbb{N}$, the connected components of the graph $G^{n},0]$ are all of type $\mathcal{F/F}$.
@@
In Section 4, it will be shown that, for all $n$, there is no infinite path in $G^{n},0]$. This fact demonstrates that, for each $n$, all the clusters of $G^{n},0]$ are finite. Note that the cycle indicated in the properties of the $\mathcal{F/F}$ components (see Theorem 3.10 in [Baccelli2017]) in each cluster of $f^{n}$-graph is the cycle of mutual nearest neighbor points, which is a cycle of length two a.s. ∎
Remark 2.
@@
It is known that there exists a directed graph $G^{\prime}$ with vertex set $\Phi^{0}$, that is constructed in a deterministic and isometry-invariant way, which is a.s. a doubly infinite path (see [Matching]). Consequently, a point-shift on $\Phi^{0}$ can be obtained by mapping each vertex to its subsequent vertex following the direction of this path. This point-shift is a bijection. By applying Mecke’s Point Stationary Theorem (see [haji_mirsadeghi_baccelli2018]), it can be concluded that the distribution of $[G,0]$ is identical to the distribution of $[G,v]$. ∎
@@
The connected components of the $\mathrm{CHN}^{2}$ EFF, $[G,0]$, belong to the $\mathcal{I/I}$ class of the foil classification theorem. In particular, under the PPP, each connected component is a one-ended tree and all of its foils (equivalence classes induced by the point-shift) are infinite a.s.
@@
Remark 3 (Evaporation and $\mathcal{I/I}$ Property).
Corollary 2.10 can also be proved using the evaporation property of the $\mathrm{CHN}^{2}$ point-shift, $f^{\infty}$, on the $\mathrm{CHN}^{2}$ EFF, $[G,0]$. Recall that a point-shift $f$ on a connected component $C\subseteq[G,0]$ evaporates $C$ if, almost surely, there is no point $x\in C$ with arbitrarily large backward iterates under $f$, i.e., $f^{-k}(x)=\emptyset$ almost surely for all sufficiently large $k$ (see [haji_mirsadeghi_baccelli2018] for more details). Corollary 24 in [haji_mirsadeghi_baccelli2018] shows that if a point-shift on a point process evaporates a connected component $C$ of its point-shift graph, then $C$ belongs to the $\mathcal{I/I}$ class. Since the intensity of exit points at level $k$ goes to zero as $k\to\infty$, it follows that $f^{\infty}$ evaporates every connected component of $[G,0]$. Hence, all such components belong to the $\mathcal{I/I}$ class.
@@
Theorem 2.11 (Connectivity of the limiting graph in dimension $d=1$).
In dimension $d=1$, under the Palm probability of $\Phi^{0}\subseteq\mathbb{R}$, the limiting $\mathrm{CHN}^{2}$ EFF is almost surely connected.
@@
By Corollary 2.7, the limiting $\mathrm{CHN}^{2}$ EFF $[G,0]$ is unimodular (as a local weak limit of the unimodular graphs $G^{n},0]$; see also [Baccelli2017]). By Corollary 2.10, each connected component of $[G,0]$ is infinite almost surely.
Assume for the sake of contradiction that $[G,0]$ is not connected. Because the construction takes place on the line and each component, at every level, connects only to one of its neighbors on the left or on the right, each connected component occupies a consecutive block of vertices along the line (with no interlacing between components). Hence, distinct components occupy disjoint order-convex blocks along $\mathbb{R}$; since every component is infinite, there can be at most two such blocks—one extending to the left and one to the right.
Let $S$ be the set containing the two extremal vertices separating these components—namely, the rightmost vertex of the left component and the leftmost vertex of the right component. These two vertices are the only vertices that have degree one in the graph and can be defined in a covariant way.
Hence, $S$ is a covariant subset of vertices in the sense of Definition 2.7 in [Baccelli2017]. Furthermore, the decomposition of $[G,0]$ into its connected components defines a covariant partition (Definition 2.9 in [Baccelli2017]). However, Lemma 2.11 (No Infinite/Finite Inclusion) in [Baccelli2017] states that, in a unimodular network, a covariant subset cannot have a finite, non-empty intersection with an infinite element of a covariant partition. Here, $S$ intersects each infinite component of the partition in exactly one vertex, violating this property. Equivalently, by Corollary 2.12 in [Baccelli2017], any covariant subset in an infinite unimodular network is almost surely either empty or infinite.Since $S$ contains exactly two vertices, this leads to a contradiction. Therefore, the limiting $\mathrm{CHN}^{2}$ EFF must be connected almost surely, which completes the proof. ∎
@@ 2.2.2 The Cardinality of Descendant Trees
More precisely, let the intensity of the cluster heads of level $k$, $\Phi_{c}^{k}$, be $\rho_{k}$. Then the mean of the cardinality of the descendant tree of a point belongs to $\Phi_{c}^{k}$, is $1/\rho_{k}$. It is known from Remark 1 that $\rho_{k}$ tend to zero, as $k$ goes to infinity.
@@
Let $[G,0]$ be the $\mathrm{CHN}^{2}$ EFF under the Palm probability of $\Phi^{0}\subseteq\mathbb{R}^{d}$. Let $N^{0}$ denote the cardinality of the descendant tree of the root 0. Then $\mathbb{E}[N^{0}]=\infty$. In particular, $N^{0}$ has a heavy tail.
@@
Since $[G,0]$ is unimodular, the Mass Transport Principle applies (see [Aldous]). Consider the $\mathrm{CHN}^{2}$ point-shift $f$ oriented from each vertex to its parent. Because every component is one-ended, the parent pointers form an infinite ancestral ray almost surely. Applying the MTP to the transport that sends one unit of mass from each vertex to all its ancestors yields $\mathbb{E}[N^{0}]=\infty$. ∎
Remark 4.
@@
3 Clustroid Hierarchical Nearest Neighbor Chain Algorithm on Cox Point Processes